# CHARME: A chain-based reinforcement learning approach for the minor embedding problem

### v1

@@ 3.4.2. Order Refining
We observe that an embedding order is a permutation of $V_{P}$, so the space of embedding orders for $P$ is equivalent to the group of permutations of $V_{P}$. Therefore, finding embedding orders with low efficiency score from this space is challenging. To address this issue, first, we establish an estimated lower bound on the efficiency scores for a special family of embedding orders. Then, based on the lower bound, we propose the subroutine Order Refining for fast exploring orders, described in Algorithm 3.
Before going further in details, we introduce notions used in this section. Given a set $S$, we define $\mathcal{E}^{(S)}$ as the group of permutations of $S$. From this, we infer that the space for embedding orders of $P$ is $\mathcal{E}^{(V_{P})}$. We also refer to an order $O\in\mathcal{E}^{(V_{P})}$ as a complete order. Considering two disjoint sets $S_{1}$ and $S_{2}$, we define the concatenation operator on the two groups $\mathcal{E}^{(S_{1})}$ and $\mathcal{E}^{(S_{2})}$ as the symbol $\oplus$. Specifically, given two orders $O_{A}=(\bar{a}_{1},\dots,\bar{a}_{|S_{1}|})\in\mathcal{E}^{(S_{1})}$ and $O_{B}=(\bar{b}_{1},\dots,\bar{b}_{|S_{2}|})\in\mathcal{E}^{(S_{2})}$, we have $O_{A}\oplus O_{B}=(\bar{a}_{1},\dots,\bar{a}_{|S_{1}|},\bar{b}_{1},\dots,\bar{b}_{|S_{2}|})\in\mathcal{E}^{(S_{1}\cup S_{2})}$. In addition, considering an order $O_{A}=\{\bar{a}_{1},\dots,\bar{a}_{|S_{1}|}\}\in\mathcal{E}^{(S_{1})}$, we denote the group of complete orders with prefix $O_{A}$ as $\mathcal{E}^{(V_{P})}_{O_{A}}\subseteq\mathcal{E}^{(V_{P})}$. Specifically, an order $O\in\mathcal{E}^{(V_{P})}_{O_{A}}$ is formed as $O=O_{A}\oplus O_{B}$ where $O_{B}\in\mathcal{E}^{(V_{P}\setminus O_{A})}$.
Next, we consider a special prefix type, named non-expansion prefix. Given a subset $S_{1}\in V_{P}$ with $t=|S_{1}|+1$, and a prefix $O_{A}=\{\bar{a}_{1},\dots,\bar{a}_{t-1}\}\in\mathcal{E}^{(S_{1})}$, we say that $O_{A}$ is a non-expansion prefix if all orders $O=(\bar{a}_{1},\dots,\bar{a}_{t-1},\bar{a}_{t},\dots,\bar{a}_{T})\in\mathcal{E}^{(V_{P})}_{O_{A}}$, results in a set of embeddings $\{\phi^{(1)},\dots,\phi^{(T)}\}$ such that with $\forall t^{\prime}\geq t$, $\phi^{(t^{\prime})}$ is non-expansion embedding.
Now, we consider constructing the embedding order by iteratively selecting an action as a node in $V_{P}$ and appending its to the current order. Each step for action selection is denoted as an embedding step. At the embedding step $t$, we assume that the embedding order $O_{A}=(\bar{a}_{1},\dots,\bar{a}_{t-1})$ is a non-expansion prefix. In other word, the topology is no longer expanded from the embedding step $t$. Theorem 1 gives us the comparison between the number of qubits gained when selecting an action $v\in V_{P}$ at the embedding steps $t$ and $t^{\prime}>t$ in this setting.
@@
Given a hardware graph $H=(V_{H},E_{H})$ and a logical graph $P=(V_{P},E_{P})$ with $T=|V_{P}|$, assume that at the embedding step $t$, we have the sequence of selected actions $O_{A}=(\bar{a}_{1},\dots,\bar{a}_{t-1})$ as a non-expansion prefix. We prove that given a node $v\in V_{P},v\notin O_{A}$, the number of qubits required to embed $v$ at the step $t$ is not greater than that required when embedding $v$ at the step $t^{\prime}$.
@@
First, we set up notions we use in this proof. We denote the embedding at the embedding step $t-1$ as $\phi^{(t-1)}:V_{P}\rightarrow V_{H}$. Specifically, the embedding $\phi^{(t-1)}$ is constructed by embedding actions $\bar{a}_{1},\dots,\bar{a}_{t-1}$ with each action as a node in the logical graph $P$. At the embedding step $t$, we denote the set of nodes in $V_{P}$ which are already embedded into the hardware as $U^{(t)}_{P}=\{v|v\in V_{P},\phi^{(t-1)}(v)\neq\emptyset\}$, while we denote the set of nodes in $V_{H}$ which are embedded by a node in the logical graph as $U^{(t)}_{H}=\{u|u\in V_{H},\exists v\in U^{(t)}_{P},u\in\phi^{(t-1)}(v)\}$. We note that the set $U^{(t)}_{P}$ is equivalent to $\{\bar{a}_{1},\dots,\bar{a}_{t-1}\}$.
We define a path in the hardware graph as $Z=\{z_{1},\dots,z_{L}\}$ where $L$ is the path length. We consider a path $Z=\{z_{1},\dots,z_{L}\}$ as a clean path at the time $t$ if $z_{i}\in V_{H}\setminus U^{(t)}_{H}$ with $i\in[1,L-1]$ and $z_{i}\in U^{(t)}_{H}$ with $i=L$. From that, we denote the set of all possible clean paths in the embedding step $t$ as $\mathcal{Z}^{(t)}$. Given a hardware node $u\in V_{H}$ and an logical node $v\in V_{P}$, we denote the set of all clean path from $u$ to the chain $\phi^{(t-1)}(v)$ as $\mathcal{Z}^{(t)}_{uv}=\{Z=\{z_{1},\dots,z_{L}\}|Z\in\mathcal{Z}^{(t)},z_{1}=u,z_{L}\in\phi^{(t-1)}(v)\}$. Note that $\mathcal{Z}^{(t)}_{uv}=\emptyset$ if $\phi^{(t-1)}(v)=\emptyset$. Then, we denote the shortest clean path at the embedding step $t$ from $u$ to $\phi^{(t-1)}(v)$ as
@@
At the embedding step $t$, given an unembedded logical node $v\in V_{P}\setminus U^{(t)}_{P}$, we denote the number of qubits gained for embedding $v$ into $H$ as $C(v|t)$. From Algorithm 1, we find the an unembedded hardware node $u\in V_{H}\setminus U^{(t)}_{H}$ such that the size of the union of shortest clean paths from $u$ to chains of neighbors of $v$ is minimized. Thus, denoted the set of neighbors of $v$ in the logical graph as $N(v)$, we can calculate $C(v|t)$ as follows:
(14) $C(v|t)=\min_{u^{\prime}\in V_{H}\setminus U^{(t)}_{H}}\left|\bigcup_{v^{\prime}\in N(v)}\tilde{Z}^{(t)}_{u^{\prime}v^{\prime}}\right|$
Next, we consider embedding the node $v$ the embedding step $t^{\prime}=t+\delta$. We assume that an arbitrary set of actions $\Delta=\{\bar{a}_{t},\dots,\bar{a}_{t+\delta-1}\}$ is embedded in intermediate embedding steps such that $v\notin\Delta$. We update the set $U^{(t^{\prime})}_{P}=U^{(t)}_{P}\cup\Delta$ and $U^{(t^{\prime})}_{H}=U^{(t)}_{H}\cup\bigcup_{i\in[t,t+\Delta-1]}\phi^{(i)}(\bar{a}_{i})$ based on the assumption that there is no expansion embedding between the embedding step $t$ and $t^{\prime}$. Then, we establish the relation between $C(v|t)$ and $C(v|t^{\prime})$. First, we observe that $U^{(t)}_{H}\subset U^{(t^{\prime})}_{H}$. Thus, we imply that $\mathcal{Z}^{(t)}\subset\mathcal{Z}^{(t^{\prime})}$ following by $\mathcal{Z}^{(t)}_{uv}\subseteq\mathcal{Z}^{(t^{\prime})}_{uv}$ for $u\in V_{H},v\in V_{P}$. As a result, from Equation 13, we have $|\tilde{Z}^{(t)}_{uv}|\leq|\tilde{Z}^{(t^{\prime})}_{uv}|$. Then, based on Equation 14, given an unembedded node $v\in V_{P}\setminus U^{(t)}_{P}$, we have:
@@
Therefore, the theorem is proven.
Input: The training logical graph $P=(V_{P},E_{P})$, hardware graph $H=(V_{H},E_{H})$, the embedding step $t$, the prefix $O_{A}=(\bar{a}_{1},\dots\bar{a}_{t-1})$, the number of recursive steps $k$, and the qubit threshold $\zeta$
Output: Return a message variable $msg$, the selected suffix $O_{B}=(\bar{a}_{t},\dots\bar{a}_{T})$, and the number of remaining recursive steps $k^{\prime}$
@@
3Initialize $\phi^{(0)}$ as an empty embedding
4for $i:=1\textrm{ to }t-1$ do
5       $P,H,\phi^{(i)}\leftarrow STATE\_TRANSITION((P,H,\phi^{(i-1)}),\bar{a}_{i-1})$
6
7if $t>T$ then
8       if $|\phi^{(T)}|<\zeta$ then
9             return True, $\varnothing$, $k-1$
@@ 5       $P,H,\phi^{(i)}\leftarrow STATE\_TRANSITION((P,H,\phi^{(i-1)}),\bar{a}_{i-1})$
11             return False, $\varnothing$, $k-1$
12
13
@@
16for $v\in V_{P}\setminus U^{(t)}_{P}$ do
17       $\phi^{\prime}:=STATE\_TRANSITION((P,H,\phi^{(t)}),v)$
@@
20
21if $\bar{F}>\zeta$ then
22       return False, $\varnothing$, $k-1$
23
@@ 26       Randomly select $v\in V_{P}\setminus U^{(t)}_{P}$
27      $msg,O_{B},k^{\prime}\leftarrow ORDER\_REFINING\left(P,H,t+1,O_{A}\oplus(v),k^{\prime}-1,\zeta\right)$
28      if $msg$ is True then
@@
30
31
32return False, $\varnothing$, 00
@@
Based on Theorem 1, given a non-expansion prefix $O_{A}=(\bar{a}_{1},\dots,\bar{a}_{t-1})$, we establish the lower bound $\bar{F}$ on the efficiency score of all embedding orders $O=(\bar{a}_{1},\dots,\bar{a}_{t-1},\bar{a}_{t},\dots,\bar{a}_{T})\in\mathcal{E}^{(V_{P})}_{O_{A}}$ as follows:
@@
Later, in the experiment, we show that for general prefix $O_{A}$, Theorem 1 is still valid for most of pairs $t$ and $t^{\prime}$, proving the usefulness of the lower bound $\bar{F}$ in general.
Then, we propose a subroutine for exploring embedding orders, described in Algorithm 3. Algorithm 3 takes a training logical graph $P=(V_{P},E_{P})$ with $T=|V_{P}|$, the hardware graph $H=(V_{H},E_{H})$, the current embedding step $t$, the prefix order $O_{A}=\{\bar{a}_{1},\dots,\bar{a}_{t}\}$, the number of recursive steps $k$ and the qubit threshold $\zeta$. Its goal is to find an suffix $O_{B}$ such that the efficiency score of $O=O_{A}\oplus O_{B}$ is less than $\zeta$ within $k$ recursive steps. Specifically, Algorithm 3 returns a message variable $msg$, the suffix $O_{B}$ and the number of remaining recursive steps $k^{\prime}$. The message variable $msg$ indicates whether a feasible suffix $O_{B}$ is found. If $msg$ is equal to $False$, $O_{B}$ is assigned as $\varnothing$.
Here, we provide a detailed explanation of Algorithm 3. First, we reconstruct embeddings $\phi^{(t-1)}$ based on Algorithm 1 (line 1 - 4). If there is no more node to consider ($t>T$), Algorithm 3 compares the efficiency score of the complete embedding order with the threshold $\zeta$ to return appropriate message variable $msg$ (line 5 - 9). Then, we calculate the estimated lower bound $\bar{F}$ based on Equation 15 (line 10 - 15). If $\bar{F}>\zeta$, Algorithm 3 stops the current recursion (line 16 - 17). Finally, a random node $v$ is picked and the next recursion at the embedding step $t+1$ is triggered with the updated prefix $O_{A}\oplus(v)$ and the number of remaining recursive steps $k^{\prime}$ (line 19 - 23). The subroutine, which is triggered at the embedding step 00, can return a complete embedding order with the efficiency score less than $\zeta$.
@@ 4. Experiments
This section presents our experimental evaluation with two focuses: 1) confirming the efficiency of our proposed exploration strategy, and 2) comparing the performance of our end-to-end RL framework, CHARME, to three state-of-the-art methods: OCT-based (Goodrich et al., 2018), Minorminer (Cai et al., 2014), and ATOM (Ngo et al., 2023a) in terms of the number of qubits used, and the total running time.

### v2

@@ 3.4.2. Order Refining
We observe that an embedding order is a permutation of $V_{P}$, so the space of embedding orders for $P$ is equivalent to the group of permutations of $V_{P}$. Therefore, finding embedding orders with a low efficiency score in this space is challenging. To address this issue, first, we establish an estimated lower bound on the efficiency scores for a special family of embedding orders. Then, based on the lower bound, we propose the subroutine Order Refining for fast exploration of orders, described in Algorithm 3.
Before going further, we introduce the notions used in this section. Given a finite set $S$, we define $\mathcal{E}^{(S)}$ as the group of permutations of $S$. From this, we infer that the space for embedding orders of $P$ is $\mathcal{E}^{(V_{P})}$. We also refer to an order $O\in\mathcal{E}^{(V_{P})}$ as a complete order. Considering two disjoint sets $S_{1}$ and $S_{2}$, we define the concatenation operator on the two groups $\mathcal{E}^{(S_{1})}$ and $\mathcal{E}^{(S_{2})}$ as the symbol $\oplus$. Specifically, given two orders $O_{A}=(\bar{a}_{1},\dots,\bar{a}_{|S_{1}|})\in\mathcal{E}^{(S_{1})}$ and $O_{B}=(\bar{b}_{1},\dots,\bar{b}_{|S_{2}|})\in\mathcal{E}^{(S_{2})}$, we have $O_{A}\oplus O_{B}=(\bar{a}_{1},\dots,\bar{a}_{|S_{1}|},\bar{b}_{1},\dots,\bar{b}_{|S_{2}|})\in\mathcal{E}^{(S_{1}\cup S_{2})}$. In addition, considering an order $O_{A}=\{\bar{a}_{1},\dots,\bar{a}_{|S_{1}|}\}\in\mathcal{E}^{(S_{1})}$, we denote the group of complete orders with the prefix $O_{A}$ as $\mathcal{E}^{(V_{P})}_{O_{A}}\subseteq\mathcal{E}^{(V_{P})}$. Specifically, an order $O\in\mathcal{E}^{(V_{P})}_{O_{A}}$ is formed as $O=O_{A}\oplus O_{B}$ where $O_{B}\in\mathcal{E}^{(V_{P}\setminus O_{A})}$.
Next, we consider a special prefix type, named a non-expansion prefix. Given a subset $S_{1}\in V_{P}$ with $t=|S_{1}|+1$, and a prefix $O_{A}=\{\bar{a}_{1},\dots,\bar{a}_{t-1}\}\in\mathcal{E}^{(S_{1})}$, we say that $O_{A}$ is a non-expansion prefix if all orders $O=(\bar{a}_{1},\dots,\bar{a}_{t-1},\bar{a}_{t},\dots,\bar{a}_{T})\in\mathcal{E}^{(V_{P})}_{O_{A}}$, result in a set of embeddings $\{\phi^{(1)},\dots,\phi^{(T)}\}$ such that with $\forall t^{\prime}\geq t$, $\phi^{(t^{\prime})}$ is a non-expansion embedding.
Now, we consider constructing the embedding order by iteratively selecting an action as a node in $V_{P}$ and appending it to the current order. Each step in action selection is denoted as an embedding step. At the embedding step $t$, we assume that the embedding order $O_{A}=(\bar{a}_{1},\dots,\bar{a}_{t-1})$ is a non-expansion prefix. In other words, the topology is no longer expanded from the embedding step $t$. Theorem 1 gives us the comparison between the number of qubits gained when selecting an action $v\in V_{P}$ at the embedding steps $t$ and $t^{\prime}>t$ in this setting.
@@
Given a hardware graph $H=(V_{H},E_{H})$, a logical graph $P=(V_{P},E_{P})$, and the step limit $T=|V_{P}|$, assume that at the embedding step $t<T$, we have the sequence of selected actions $O_{A}=(\bar{a}_{1},\dots,\bar{a}_{t-1})$ as a non-expansion prefix. Given a node $v\in V_{P},v\notin O_{A}$, the number of qubits required to embed $v$ at step $t$ is not greater than that required when embedding $v$ at step $t^{\prime}$ with $t<t^{\prime}\leq T$.
@@
We denote the embedding at the embedding step t1t-1 as $\phi^{(t-1)}:V_{P}\rightarrow V_{H}$. Specifically, the embedding $\phi^{(t-1)}$ is constructed by embedding actions $\bar{a}_{1},\dots,\bar{a}_{t-1}$ with each action as a node in the logical graph $P$. At the embedding step $t$, we denote the set of nodes in $V_{P}$ that are already embedded in the hardware as $U^{(t)}_{P}=\{v|v\in V_{P},\phi^{(t-1)}(v)\neq\emptyset\}$, while we denote the set of nodes in $V_{H}$ that are embedded by a node in the logical graph as $U^{(t)}_{H}=\{u|u\in V_{H},\exists v\in U^{(t)}_{P},u\in\phi^{(t-1)}(v)\}$. We note that the set $U^{(t)}_{P}$ can be considered as an unordered representation of the sequence $O_{A}=\{\bar{a}_{1},\dots,\bar{a}_{t-1}\}$ because it contains all elements of $O_{A}$. For mathematical convenience in this proof, we primarily use the set $U^{(t)}_{P}$ to represent the sequence $O_{A}$.
We define a path of length $L$ in the hardware graph as $Z=\{z_{1},\dots,z_{L}\}$ with $(z_{i},z_{i+1})\in E_{H},i\in[1,L-1]$. We consider a path $Z=\{z_{1},\dots,z_{L}\}$ to be a clean path at time $t$ if $z_{i}\in V_{H}\setminus U^{(t)}_{H}$ with $i\in[1,L-1]$ and $z_{i}\in U^{(t)}_{H}$ with $i=L$. From that, we denote the set of all possible clean paths in the embedding step $t$ as $\mathcal{Z}^{(t)}$. Given a hardware node $u\in V_{H}$ and a logical node $v\in V_{P}$, we denote the set of all clean paths from $u$ to the chain $\phi^{(t-1)}(v)$ as $\mathcal{Z}^{(t)}_{uv}=\{Z=\{z_{1},\dots,z_{L}\}|Z\in\mathcal{Z}^{(t)},z_{1}=u,z_{L}\in\phi^{(t-1)}(v)\}$. Note that $\mathcal{Z}^{(t)}_{uv}$$\emptyset$ if $\phi^{(t-1)}(v)=\emptyset$. Then, we denote the shortest clean path at embedding step $t$ from $u$ to $\phi^{(t-1)}(v)$ as
@@
At embedding step $t$, given an unembedded logical node $v\in V_{P}\setminus U^{(t)}_{P}$, we denote the number of qubits gained for embedding $v$ into $H$ as $C(v|t)$. From Algorithm 1, we find an unembedded hardware node $u\in V_{H}\setminus U^{(t)}_{H}$ such that the size of the union of the shortest clean paths from $u$ to chains of neighbors of $v$ is minimized. Thus, denoting the set of neighbors of $v$ in the logical graph as $N(v)$, we can calculate $C(v|t)$ as follows:
[…]
@@
Therefore, the theorem is proven. ∎
Input: The training logical graph $P=(V_{P},E_{P})$, hardware graph $H=(V_{H},E_{H})$, the embedding step $t$, the current embedding $\phi$, the prefix $O_{A}=(\bar{a}_{1},\dots\bar{a}_{t-1})$, the number of recursive steps $k$, and the qubit threshold $\zeta$
Output: Return a message variable msgmsg, the selected suffix $O_{B}=(\bar{a}_{t},\dots\bar{a}_{T})$, and the number of remaining recursive steps $k^{\prime}$
@@
3if $t>T$ then
4 if $|\phi|<\zeta$ then
5    return True, $\varnothing$, k1k-1
@@
7    return False, $\varnothing$, k1k-1
8
9
@@
12for $v\in V_{P}\setminus U^{(t)}_{P}$ do
13 $\phi^{\prime}:=\texttt{State Transition}((P,H,\phi),v)$
@@
16
17if $\bar{F}>\zeta$ then
18 return False, $\varnothing$, k1k-1
19
@@ 22   Randomly select $v\in V_{P}\setminus U^{(t)}_{P}$
23 $\phi^{\prime}:=\texttt{State Transition}((P,H,\phi),v)$
24 $msg,O_{B},k^{\prime}\leftarrow\texttt{Order Refining}\left(P,H,t+1,\phi^{\prime},O_{A}\oplus(v),k^{\prime}-1,\zeta\right)$
25 if msgmsg is True then
@@ 23 $\phi^{\prime}:=\texttt{State Transition}((P,H,\phi),v)$
27
28
29return False, $\varnothing$, 0
@@
Based on Theorem 1, given a non-expansion prefix $O_{A}=(\bar{a}_{1},\dots,\bar{a}_{t-1})$, we establish the lower bound $\bar{F}$ on the efficiency score of all embedding orders $O=(\bar{a}_{1},\dots,\bar{a}_{t-1},\bar{a}_{t},\dots,\bar{a}_{T})\in\mathcal{E}^{(V_{P})}_{O_{A}}$ as:
@@
Subsequently, in our experiments, we show that for the general prefix $O_{A}$, Theorem 1 is still valid for most pairs $t$ and $t^{\prime}$, proving the usefulness of the lower bound $\bar{F}$ in general.
We now introduce a subroutine to explore embedding orders, described in Algorithm 3. Algorithm 3 takes a training logical graph $P=(V_{P},E_{P})$ with $T=|V_{P}|$, the hardware graph $H=(V_{H},E_{H})$, the current embedding step $t$, the current embedding $\phi$, the prefix order $O_{A}=\{\bar{a}_{1},\dots,\bar{a}_{t-1}\}$, the number of recursive steps $k$, and the qubit threshold $\zeta$. It is important to note that $\phi$ is the embedding of nodes in the prefix $O_{A}$. The objective is to find a suffix $O_{B}$ such that the efficiency score of $O=O_{A}\oplus O_{B}$ is less than $\zeta$ within $k$ recursive steps. Specifically, Algorithm 3 returns a message variable msgmsg, the suffix $O_{B}$, and the number of remaining recursive steps $k^{\prime}$. The message variable msgmsg indicates whether a feasible suffix $O_{B}$ is found. If msgmsg is equal to $False$, $O_{B}$ is assigned as $\varnothing$.
Here, we provide a detailed explanation of Algorithm 3. First, we set $T$ as the total number of nodes in the logical graph (line 1). If there are no more nodes to consider ($t>T$), Algorithm 3 compares the qubit usage of the current embedding $\phi$ with the threshold $\zeta$, returning the appropriate message variable msgmsg (lines 2 - 6). Then, the estimated lower bound $\bar{F}$ is calculated based on Equation 17 (lines 7 - 12). If $\bar{F}$ exceeds the qubit threshold $\zeta$, Algorithm 3 terminates the current recursion (lines 13 - 14). Subsequently, a random node $v$ is selected and the next recursion at the embedding step $t+1$ is initiated with the updated prefix $O_{A}\oplus(v)$ and the number of remaining recursive steps $k^{\prime}$ (lines 16 - 21). The subroutine, triggered at embedding step 0, can return a complete embedding order with an efficiency score less than $\zeta$.
@@ 4. Experiments
This section presents our experimental evaluation with two focuses: 1) confirming the efficiency of our proposed exploration strategy, 2) showing the limitation of NaiveRL and 3) showing the superiority of CHARME by comparing its performance to three state-of-the-art methods: OCT-based (Goodrich et al., 2018),Minorminer (Cai et al., 2014), and ATOM (Ngo et al., 2023a) in terms of the number of qubits used and the total running time.